OpenStudy (anonymous):

what is the sum of the arithmetic sequence 3 9 15... if there are 36 terms

OpenStudy (anonymous):

first we will find the common difference by subtracting the first term from the second term. 9 - 3 = 6 so our common difference (d) is 6 Now we will find the 36th term because we will need it to find the sum an = a1 + (n - 1) * d n = the term we want to find = 36 a1 = first term = 3 d = common difference = 6 now we sub a36 = 3 + (36 - 1) * 6 a36 = 3 + 35 * 6 a36 = 3 + 210 a36 = 213 now that we know the 36th term. we can find the sum sn = n(a1 + an) / 2 sn = sum of n terms = 36 a1 = first term = 3 an = 36th term = 213 now we sub s36 = 36(3 + 213) / 2 s36 = 36 (216) / 2 s36 = 7776/2 s36 = 3888 If I am not mistaken, the sum of the 36 terms is 3888